OpenStudy (anonymous):

Need help 10th grade geometry... Two realativly small aquafers are the Rush Springs (RS) Aquafer and the Arbuckle Simpson (AS) Aquafer, both in Oklahoma. Suppose that on a certain day in 1985, 52 million gallons of water perday were taken from RS aquafer, and 8 million gallons of water per day were taken from the AS aquafer. If the RS aquifer began with 4500 million gallons of was and AS aquafer began with 3000 million gallons of water and no rain fell, write a slope-intercept equation for each aquafer and find out how manyh days passed untill both aquifers held the same about of water. (ro

OpenStudy (anonymous):

RS = 4500 - 52t AS = 3000 - 8t They'll have the same amount of water when they're equal: 4500 - 52t = 3000 - 8t 44t = 1500 t = 34.09 days.